# SeagrassFinder: Deep Learning for Eelgrass Detection and Coverage Estimation in the Wild

### v1

@@ III-A Initial Model Evaluation
The results of our initial model evaluation experiments are shown in Tbl. IV. Based on the high level of AUROC scores, and a low variation for all models, we ascertain that the task of eelgrass classification can in fact be conducted by DNN models. Secondly, the models trained and tested on UW-enhanced images show a higher accuracy compared to their non-enhanced counterparts in 26/30 cases. Thus UW-enhancement has a positive effect the models’ ability to predict eelgrass presence accurately in the context of transfer learning with pre-training on Imagenet. We conclude that some models may benefit from UW-image enhancement. Generally we find that the Vision Transformer and DenseNet architecture show good performance on the target task. Surprisingly the smallest ResNet model, ResNet-18++, gives the best results across all three metrics among the ResNet models. The Vision Transformer shows the best results overall compared to all DNN architectures across all three metrics.
@@
TABLE IV: Test results from initial experiments, with models trained on LYT-5, LYT-10, LYT-12 and LYT-14 and tested on LYT-20. The 3 metrics of interest are test accuracy, are under ROC curve (AUROC) and test calibration error (CE). Higher scores mean better performance and are marked by ($\uparrow$) and vice versa marked by ($\downarrow$). Models trained and tested on UW-enhanced images are marked by ++.
@@
III-B Leave-One-Transect-Out Results
As can be seen in Tbl. V, the ViT architecture continuously shows the best performance, with the enhanced version ViT++ performing best. Using models on UW-enhanced data also continues to provide the highest model scores across all model architectures. We also notice that the smaller DenseNet-169++ model has a higher accuracy than the DenseNet-201, although performing worse in the AUROC and CE categories. This highlights the importance of using multiple different metrics to evaluate models.
TABLE V: Leave-one-transect-out cross validation results. The 3 metrics of interest are test accuracy, area under ROC curve (AUROC) and test calibration error (CE). Higher scores mean better performance and are marked by ($\uparrow$) and vice versa marked by ($\downarrow$). Models trained and tested on UW-enhanced images are marked by ++.
@@ III-B Leave-One-Transect-Out Results
III-C Final Test Dataset Model Results
We report our final model test results in Tbl. VI. In the final model training and testing, for non-enhanced models both DenseNet-201 and Vision Transformer achieve high scores, with both models gaining high AUROC scores on the test set. Both models also achieved very low calibration error scores, 0.084 and 0.086 respectively, showing both models capability to not only predict accurately, but also confidently. Overall the Vision Transformer performed best across all metrics. Both models showed an increase in performance when trained and tested on the UW-enhanced images. With a total training time of around 30 minutes we show that our transfer learning approach is highly efficient, compared to training a ViT from scratch which can take up to three days on four GPUs (Touvron et al., 2021).
TABLE VI: Final Model Results: Models trained and tested on enhanced images are marked by ++.
@@ III-C Final Test Dataset Model Results
III-D Eelgrass Presence and Coverage Estimation
@@
We map the ViT predictions from LYT-9, LYT-10, LYT-11, LYT-13, LYT-19, LYT-20 geographically in Fig. 8(a). The green marks denoting the presence of eelgrass, and yellow the absence. In Fig. 8(b) all eelgrass present data points are mapped on a heat map, giving us a good visual estimation for eelgrass coverage in the area. We also present information on the extracted eelgrass distribution relative to water depth. This relationship is plotted relative to a normalized water depth bin, since the water depth data is not equally distributed (see Fig. 9).
Since we do not have expert labels on eelgrass coverage available of the LYT transects, we show how our coverage estimation compares with manual labels of a different EIA transect video. In Fig. 10 we show the eelgrass coverage estimation based on our ViT predictions compared to video analysis labeled by experts. The respective transect video was taken at a location around 150km away from the LYT video transects. We can see that our estimation based on temporal means of ViT predictions shows spikes in similar timestamps compared to the expert labels. Figure 10 demonstrates that our method is generalizable to an environment with a significantly different seabed.
@@
Figure 8: (a) Eelgrass presence: the lines mark transect surveys conducted in the Copenhagen harbor along planned survey paths, green dots indicate detected eelgrass presence along each transect, yellow dots indicate eelgrass absence. (b) Eelgrass coverage: an estimation of eelgrass coverage in the benthic environment represented by a heat map of the transects.
Figure 9: Eelgrass presence percentage categorized by depth (m).
@@
Figure 10: Visual Coverage Estimation: (blue) our new approach based on the temporal mean of ViT predictions of a transect video, (orange) human annotated eelgrass coverage values.
IV Discussion
IV-A Video Data Limitations and Solutions
@@
During the sailing of the transect surveys, the recorded video footage is streamed to video processing software which adds GPS location, water depth and heading (i.e. course over ground measured in degrees) to each video as an overlay. Streaming lag can cause a delay in the timestamps in the imprinted overlay. This is problematic when trying to align the images to GPS locations during post-processing. Additionally there is a mismatch ($\sim$ 5-20m) between the imprinted GPS locations of the vessel and the actual camera position on the towed sled. One approach to improve GPS position accuracy, and generally make steps towards automated eelgrass inspection is to deploy Autonomous Underwater Vehicles (AUVs). AUVs have multiple benefits, for example more accurate position estimation compared to a towed sled, which can be achieved by fusing information from additional location sensors. This work is directly applicable to a scenario where an AUV with a forward-facing camera is deployed to autonomously perform EIAs without a manual operator. The AUV might have a high-level objective of inspecting seagrass meadows and therefore needs to detect the meadow boundaries in real-time, as for example demonstrated by Ruscio et al. (Ruscio et al., 2023). With our analysis on DNN performance, a robotics engineer can choose a suitable model according to hardware constraints for integration into the AUV navigation stack.
Finally, during inspection of some of the transect images, we noticed motion blur in some frames. This was initially attributed to the camera’s low quality and frame-rate and later suspected to stem from an automated video compression. To overcome this we recommend investing in a high-quality waterproof camera system. It is important to mention that humans are good at filtering out such motion trails, while DNN models are not, due to their dependence on raw pixel data. Therefore this should be taken into consideration when using this dataset for future applications.
IV-B Eelgrass Coverage Estimation
@@
In this work, we present a novel idea for eelgrass coverage estimation to overcome practical constraints of traditional methods. We outline the benefits and limitations of our approach in Sect. II-F. Additionally we argue, that human labeling inherently suffers from observer bias. Although DNNs are not immune to biases, they can reduce variability across images and perspectives as well as offer a consistent and deterministic analytical process, once they are trained. Human labeling also suffers from the tendency of annotators to focus disproportionately on specific parts of an image. This issue is particularly pronounced here since our images are captured at an angle towards the seafloor; eelgrass in the foreground might induce the annotator to indicate a high coverage percentage even if there is little eelgrass present in the background. We assert that our data-driven approach of calculating the temporal mean of our DNN-based binary predictions offers a robust and accurate approach to estimating eelgrass coverage in the wild. Here we would like to stress that this approach should act as a supportive tool for marine biologists to conduct EIAs in a more data-driven manner. This tools presented in this work should not be seen as a replacement for expert domain knowledge, rather, a showcase to apply state-of-the-art technologies to support marine experts in their work.
IV-C Underwater Image Enhancement
@@
We find that applying Deep WaveNet (Sharma et al., 2023) as an UW image enhancement tool proves to be effective in improving model accuracy. However, the computational demands are substantial, requiring approximately three hours of processing time on an A10G GPU for the entire training set, or around 1.2 seconds per image. The trade-off between the computational investment and the incremental performance gains must be carefully considered in practice, given the rising carbon footprint of training of DNNs (Luccioni and Hernandez-Garcia, 2023). Additionally, it is also important to consider the potential for the the Deep WaveNet model to introduce bias in eelgrass presence classification based on water depth. As Sharma et al. show, the receptive field size of global features is more effective at greater water depths, than in shallow waters. Since all transects used for model training and testing occur in water depths of 2-14m, the impact of Deep WaveNet in this context is minimal but cannot be entirely discounted.
V Conclusion
@@
This work shows the potential of deep learning models to automate the process of seagrass detection in underwater video transects from challenging visual conditions in the wild. By training a deep neural networks to classify underwater images into “Eelgrass Present” and “Eelgrass Absent” classes, our proposed methodology allows for the efficient processing of large volumes of video data collected in environmental impact assessments, enabling the acquisition of much more detailed information on seagrass distributions compared to current manual annotation methods. Our results show that deep learning models, particularly Vision Transformers, can achieve high performance in predicting eelgrass presence, with AUROC scores exceeding 0.95 on the final test dataset. We also show an effective implementation of transfer learning, and that the application of the underwater image enhancement model further improves model’s capabilities.
@@
VI Acknowledgments
The authors would like to thank their collaborators at DHI: Aron Lank Jensen, Jesper Goodley Dannisøe and Lindsey Aies. This project was funded and run in collaboration with DHI (DHI, 2023). This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 956200.

### v2

@@ 4.2 Initial Model Evaluation
The results of our initial model evaluation experiments are shown in Tbl. 4. Based on the high level of AUROC scores, and a low variation for all models, we ascertain that the task of eelgrass classification can in fact be conducted by DNN models. Secondly, the models trained and tested on UW-enhanced images show a higher accuracy compared to their non-enhanced counterparts in 26/30 cases. Thus UW-enhancement has a positive effect on the models’ ability to predict eelgrass presence accurately in the context of transfer learning with pre-training on Imagenet. We conclude that some models may benefit from UW-image enhancement. Generally we find that the Vision Transformer and DenseNet architecture show good performance on the target task. Surprisingly the smallest ResNet model, ResNet-18++, gives the best results across all three metrics among the ResNet models. The Vision Transformer shows the best results overall compared to all DNN architectures across all three metrics.
@@
Table 4: Test results from initial experiments, with models trained on LYT-5, LYT-10, LYT-12 and LYT-14 and tested on LYT-20. The 3 metrics of interest are test accuracy, are under ROC curve (AUROC) and test calibration error (CE). Higher scores mean better performance and are marked by ($\uparrow$) and vice versa marked by ($\downarrow$). Models trained and tested on UW-enhanced images are marked by ++.
@@
4.3 Leave-One-Transect-Out Validation
As can be seen in Tbl. 5, the ViT architecture continuously shows the best performance, with the enhanced version ViT++ performing best. Using models on UW-enhanced data also continues to provide the highest model scores across all model architectures. We also notice that the smaller DenseNet-169++ model has a higher accuracy than the DenseNet-201, although performing worse in the AUROC and CE categories. This highlights the importance of using multiple different metrics to evaluate models.
Table 5: Leave-one-transect-out cross validation results. The 3 metrics of interest are test accuracy, area under ROC curve (AUROC) and test calibration error (CE). Higher scores mean better performance and are marked by ($\uparrow$) and vice versa marked by ($\downarrow$). Models trained and tested on UW-enhanced images are marked by ++.
@@ 4.3 Leave-One-Transect-Out Validation
4.4 Final Model Evaluation
We report our final model test results in Tbl. 6. In the final model training and testing, for non-enhanced models both DenseNet-201 and Vision Transformer achieve high scores, with both models gaining high AUROC scores on the test set. Both models also achieved very low calibration error scores, 0.084 and 0.086 respectively, showing both models capability to not only predict accurately, but also confidently. Overall the Vision Transformer performed best across all metrics. Both models showed an increase in performance when trained and tested on the UW-enhanced images. With a total training time of around 30 minutes we show that our transfer learning approach is highly efficient, compared to training a ViT from scratch which can take up to three days on four GPUs (Touvron et al., 2021).
Table 6: Final Model Results: Models trained and tested on enhanced images are marked by ++.
@@ 4.4 Final Model Evaluation
4.5 Eelgrass Coverage Estimation
@@
We map the ViT predictions from LYT-9, LYT-10, LYT-11, LYT-13, LYT-19, LYT-20 geographically in Fig. 9(a). The green marks denote the presence of eelgrass, and yellow the absence. In Fig. 9(b) all eelgrass present data points are mapped onto a heat map, giving us a good visual estimation for eelgrass coverage in the area.
Since we do not have expert labels on eelgrass coverage available of the LYT transects, we show how our coverage estimation compares with manual labels of a different EIA transect video. In Fig. 10 we show the eelgrass coverage estimation based on our ViT predictions compared to video analysis labeled by experts. The respective transect video was taken at a location around 150km away from the LYT video transects. Figure 10 serves as a qualitative demonstration of how our coverage estimation method captures ecologically meaningful patterns in an automated and efficient manner. Figure 10 also shows that our method is able to generalizable to an environment with a substantially different seabed. Additionally, we computed a Spearman correlation test for the two time series which is $0.3169$ with a p-value of $2.2331\times 10^{-8}$. However it is important to note, that expert labels are sparse, highly subjective and often not repeatable, therefore this result should be interpreted with caution as there is no baseline for what constitutes a “good” value. Selecting an appropriate metric and establishing its utility for experts is an avenue for further research.
@@
Figure 9: (a) Eelgrass presence: the lines mark transect surveys conducted along planned survey paths, green dots indicate detected eelgrass presence along each transect, yellow dots indicate eelgrass absence. (b) Eelgrass coverage: an estimation of eelgrass coverage in the benthic environment represented by a heat map of the transect surveys.
@@
Figure 10: Visual Coverage Estimation: (blue) our new approach based on the temporal mean of ViT predictions of a transect video, (orange) human annotated eelgrass coverage.
5 Discussion
Here we discuss the ecological importance of our findings, as well as limitations, future work and applications.
5.1 Video Data Limitations and Solutions
@@
During the sailing of the transect surveys, the recorded video footage is streamed to video processing software which adds GPS location, water depth and heading (i.e. course over ground measured in degrees) to each video as an overlay. Streaming lag can cause a delay in the timestamps in the imprinted overlay. This is problematic when trying to align the images to GPS locations during post-processing. Additionally there is a mismatch ($\sim$ 5-20m) between the imprinted GPS locations of the vessel and the actual camera position on the towed sled. One approach to improve GPS position accuracy, and generally make steps towards automated eelgrass inspection, is to deploy more advanced multi-purpose Remotely Operated Towed Vehicles (ROTVs) or Autonomous Underwater Vehicles (AUVs). Compared to our basic mechanical towed sled, ROTVs offer many advantages; they can carry higher payloads and thus multiple onboard sensors. If fused together, the data streams of these sensors can provide more detailed and accurate insights of the benthic environment of transect surveys.
5.2 Automation of Seagrass Monitoring
The automation of seagrass detection and coverage estimation is crucial for enabling efficient, timely and consistent monitoring of the health of underwater ecosystems. By eliminating the need for manual annotation, automated methods allow for processing of high volumes of video data and thus enable us to detect and react to ecological changes in ecosystems and respond more rapidly for example through restoration efforts by transplantation (Waycott et al., 2009). Repeated and data-driven monitoring also enables the assessment of environmental impacts associated with coastal developments such as tunnels or windfarms, that is, before, during and after construction. Since seagrass meadows also act as nursery grounds for fish, it is estimated that 20% of the global fishery productivity is provided by seagrass meadows (Unsworth et al., 2019b). As such, automated and repeated seagrass monitoring can potentially serve as an indirect indicator of local fish stock health. Automated seagrass coverage estimation could also serve as a valuable addition to blue carbon strategies and aid in tracking climate mitigation targets (Unsworth et al., 2019a; Fourqurean et al., 2012). While seagrass detection in a benthic environments provides information about the presence and locality of seagrass meadows, tracking seagrass coverage offers more nuanced insights into their condition, density and health. Activities like dredging, sedimentation, or eutrophication may reduce eelgrass density before total loss and is therefore an important advancement for marine conservation and coastal management. Here we would like to emphasize the importance of drawing a clearer distinction between spatial seagrass coverage, which refers to its actual density of vegetation, and pixel-based coverage, which simply measures the proportion of image pixels identified as seagrass. To the best of our knowledge, this distinction is currently overlooked in computer vision approaches to seagrass coverage estimation.
5.3 Automated Transect Surveys with AUVs
Deploying Autonomous Underwater Vehicles (AUVs) has multiple benefits for automated transect surveying compared to a towed sled. An AUV with a forward-facing camera can perform EIAs autonomously, without the need for a manual operator. The AUV might have the high-level objective of inspecting seagrass meadows and therefore needs to detect the meadow boundaries in real-time, as for example demonstrated by Ruscio et al. (Ruscio et al., 2023). Our analysis of DNN performance enables robotics engineers to now choose a suitable DNN according to their specific hardware constraints, and integrate a trained DNN into the AUV navigation stack for an improved inspection mission protocol. Simply increasing the operating velocity of the AUV in areas devoid of seagrass can cause substantial cost reductions, particularly in the rental of research vessels. In addition to this, AUVs are able to perform more advanced missions than simple transect lines. A typical inspection mission can be conducted in form of a lawnmower pattern, which covers a much larger area and thus gives a better understanding of the health of the underwater environment.
5.4 Eelgrass Coverage Estimation
@@
In this work, we present a novel idea for eelgrass coverage estimation to overcome practical constraints of traditional methods. We outline the benefits and limitations of our approach in Sect. 2.6. Additionally we argue, that human labeling inherently suffers from observer bias. Although DNNs are not immune to biases, they can reduce variability across images and perspectives as well as offer a consistent and deterministic analytical process, once they are trained. Human labeling also suffers from the tendency of annotators to focus disproportionately on specific parts of an image. This issue is particularly pronounced here since our images are captured at an angle towards the seafloor; eelgrass in the foreground might induce the annotator to indicate a high coverage percentage even if there is little eelgrass present in the background. We assert that our data-driven approach of calculating the temporal mean of our DNN-based binary predictions offers a robust and accurate approach to estimating eelgrass coverage in the wild. Here we would like to stress that this approach should act as a supportive tool for marine biologists to conduct EIAs in a more data-driven manner. The tools presented in this work should not be seen as a replacement for expert domain knowledge, rather, a showcase to apply state-of-the-art technologies to support marine experts in their work. With our methodology, marine experts can finally focus their efforts on interpretation of environmental impact instead of data annotation. This increases quality assurance, but also employee well-being by reducing time spent on tedious manual tasks. Additionally, we suspect that our methodology and analytic setup can easily be expanded to automatically detect other taxonomically related species of submerged aquatic vegetation with similar visual characteristics.
5.5 Data Availability and Underwater Image Quality
There is a huge growth in interest for underwater image analysis, not only for EIAs, but also in AUV camera footage and static underwater cameras. However, this discipline has many hurdles. First, the limited availability of relevant and diverse data sets poses a hard problem, not only for marine biologists but also researchers in computer vision. Second, the amount of financial resources required to commission a research vessels can be extremely high. Even for static camera setups underwater, we require advanced technologies for efficiently managing battery power and protecting hardware from water damage. To overcome the sparsity of labeled data sets, Abid et al. (Abid et al., 2024) propose unsupervised curriculum learning to address uncertainties in seagrass classification. Here, the learning algorithm chooses image samples in order of difficulty, from easiest to hardest, for a more robust learning strategy on the DeepSeagrass data set (Raine et al., 2020).
The visual quality of underwater images taken in the wild poses another big challenge. Some visual aspects such as back scatter, haze effects, caustics, low light and low contrast can lead to uncertainties in the DNN predictions and potential misclassifications. In the context of transfer learning in medical images, recent work has shown that the choice of the source data for DNN pretraining has an effect on the performance of the subsequent target task (Juodelyte et al., 2024). We are curious whether selecting an alternative domain dataset to ImageNet for pretraining could mitigate the domain shift challenges observed from on-land to underwater images. This investigation may improve DNN accuracy and generalization, however this remains out of scope.
During visual inspection of the transect videos, we noticed motion blur in some images. This was initially attributed to the camera’s low quality and frame-rate and later suspected to stem from an automated video compression. To overcome this we recommend investing in a high-quality waterproof camera system. It is important to mention that humans are good at filtering out such motion trails, while DNNs are not, due to their dependence on raw pixel data. Therefore this should be taken into consideration when using this dataset for future applications.
5.6 Underwater Image Enhancement
@@
We find that applying Deep WaveNet (Sharma et al., 2023) as an UW image enhancement tool proves to be effective in improving model accuracy. However, the computational demands are substantial, requiring approximately three hours of processing time on an A10G GPU for the entire training set, or around 1.2 seconds per image. The trade-off between the computational investment and the incremental performance gains must be carefully considered in practice, given the rising carbon footprint of training of DNNs (Luccioni and Hernandez-Garcia, 2023). Additionally, it is also important to consider the potential for the Deep WaveNet model to introduce bias in eelgrass presence classification based on water depth. As Sharma et al. show, the receptive field size of global features is more effective at greater water depths, than in shallow waters. Since all transects used for model training and testing occur in water depths of 2-14m, the impact of Deep WaveNet in this context is minimal but cannot be entirely discounted.
6 Conclusion
@@
This work shows the potential of deep neural networks to automate the process of seagrass detection in underwater videos from challenging visual conditions in the wild. By training deep neural networks to classify underwater images into “Eelgrass Present” and “Eelgrass Absent” classes, our proposed methodology allows for the efficient processing of large volumes of video data collected in environmental impact assessments, enabling the acquisition of much more detailed information on seagrass distributions compared to current manual annotation methods. Our results show that deep neural networks, particularly Vision Transformers, can achieve high performance in predicting eelgrass presence, with AUROC scores exceeding 0.95 on the final test dataset. We also show an effective implementation of transfer learning, and that the application of the underwater image enhancement model further improves models’ capabilities.
@@
7 Acknowledgments
The authors would like to thank their collaborators at DHI: Aron Lank Jensen, Jesper Goodley Dannisøe and Lindsey Aies. We would like to thank “By & Havn” for kindly allowing us to use and share their data, a crucial resource for our findings and the research community. This project was funded and run in collaboration with DHI (DHI, 2023). This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 956200.
8 Data Availability
The data used to support our findings is titled “SeagrassFinder: An Underwater Eelgrass Image Classification Dataset” and is publicly available on zenodo:
https://zenodo.org/records/13904604.
[…]
@@ Appendix A Appendix
A.1 Annotation Analysis
We report Cohens kappa scores for all annotators in Fig. 11 and mistake percentage and disagreement percentage in Fig. 12.
@@ A.1 Annotation Analysis
A.2 Glossary of Acronyms
We provide a glossary with all acronyms used in our work for reference in Tbl. 7.
Table 7: A list of all acronyms used in this work in alphabetical order.
Acronym Description AP Annotation Platform AUV Autonomous Underwater Vehicle AUROC Area Under Receiver Operating Characteristic Curve CE Calibration Error CMOS Complementary Metal Oxide Semiconductor DNN Deep Neural Network EIA Environmental Impact Asessment FPS Frames Per Second GPS Global Positioning System GPU Graphical Processing Unit LYT Lynetteholm Project RGB Red, Green, Blue ROTV Remotely Operated Towed Vehicles SQL Structured Query Language TM Temporal Mean URL Uniform Resource Locator UW Underwater ViT Vision Transformer
A.3 Screenshot of SeagrassFinder AP
We provide a screenshot of the labeling instruction page of SeagrassFinder AP in Fig. 13.
Figure 13: A screenshot of the start page of the SeagrassFinder AP showing the visual guidance for the labeling instructions.
A.4 Eelgrass Presence by Water Depth
Based on the coverage estimations we can additionally extract eelgrass distribution relative to water depth, as it is an important factor in environmental impact assessments. This metric can be used as a proxy for estimating any human impact on the local marine ecosystem, as such impacts typically become most evident at the deepest limits of seagrass distribution. Transect surveys are specifically used to determine seagrass abundance over a range of water depths (Short and Coles, 2001). We present our results in Fig. 14.
Figure 14: Eelgrass presence percentage sorted by depth (m).